OpenStudy (drshinigami):

What is the equation of the line perpendicular to –x + y = 7 and passing through (-1, -1)?

OpenStudy (drshinigami):

-x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0 theses are the choices

OpenStudy (_malice_):

r.i.p queenie, death by typing a reply. 2016

OpenStudy (queenie):

-x + y = 7 y = x + 7...the slope of this line is 1 or 1/1.....a perpendicular line will have a negative reciprocal slope. All that means is flip the slope and change the sign. So the perpendicular slope we need is -1. now we use y = mx + b....(-1,-1)...x = -1 and y = -1 slope(m) = -1 now we sub and find b, the y intercept. y = mx + b -1 = -1(-1) + b -1 = 1 + b -1 - 1 = b -2 = b so ur perpendicular line will be : y = -x -2 or x + y = -2 or -x - y = 2

OpenStudy (queenie):

sorry it took so long....I couldn't get my answer to match ur answer choices....finally got it and it is -x - y = 2

OpenStudy (drshinigami):

thank you queenie

OpenStudy (queenie):

lol....yw